OpenStudy (anonymous):

Solve. 1/x-1 + 1/4x-4=5/4

OpenStudy (johnweldon1993):

\[\large \frac{1}{x - 1} + \frac{1}{4x - 4} = \frac{5}{4}\] like that?

OpenStudy (anonymous):

Yes

OpenStudy (johnweldon1993):

Notice how if you multiply the left-most fraction by 4/4 you will receive the same denominator as the right fraction. \[\large \frac{1}{x-1} \times \frac{4}{4} = \frac{4}{4x - 4}\] right? So now you have \[\large \frac{4}{4x- 4} + \frac{1}{4x - 4} = \frac{5}{4}\] What can you do from here?

OpenStudy (johnweldon1993):

We can combine from fractions over the common denominator \[\large \frac{4}{4x - 4} + \frac{1}{4x - 4} \rightarrow \frac{4 + 1}{4x - 4}\] So now we have \[\large \frac{5}{4x - 4} = \frac{5}{4}\] And we can then solve for 'x'

OpenStudy (anonymous):

How do we solve for x with this problem?

OpenStudy (johnweldon1993):

Well we can do it a long way...or a quick way... The quick way would be...to notice that we already have the 5 on top like we want... So we need to make the bottom become the 4 that we want... The bottom is 4x - 4 and we want that to = 4... so lets solve 4x - 4 = 4

OpenStudy (anonymous):

ok